OpenStudy (anonymous):

1. Solve: 34y−2=6−14y A. y = 4 B. y = 16 C. y = 2 D. y = 8

OpenStudy (anonymous):

34y−2=6−14y 34y +14y = 6+2 48y = 8 y = 8/48